OpenStudy (anonymous):

Is {-1/3,3} a solution to 3x + 8y = 23?

OpenStudy (aroub):

-1/3 , 3 3x+8y=23 3(-1/3)+8(3)=23 -1+24=23 23=23 so, yes :)